OpenStudy (firejay5):

Find the value of x to the nearest degree. Show work and explain in order to get medal for patronage! :D

OpenStudy (firejay5):



OpenStudy (anonymous):

First, you need to use Pythagoras. a^2 + b^2 = c^2

OpenStudy (anonymous):

In this case, c would be 3sqrt(5) Therefore, the equation you would use is (3(sqrt(5))^2 - 3^2 = b^2.

OpenStudy (anonymous):

This would give you a b value of 6. Now once you have six, you can use the Primary Trig Ratios to find x.

OpenStudy (anonymous):

Wait, you only need x?

OpenStudy (anonymous):

In that case, all you need to do is use the Primary Trig Ratios to start out with. cosx = adjacent/hypotenuse = 3/(3sqrt(5))

OpenStudy (anonymous):

Therefore, x = cos^-1 (3/(3sqrt(5))

OpenStudy (anonymous):

\[\cos x=\frac{ 3 }{ 3\sqrt{5} }=\frac{ 1 }{ \sqrt{5} }\] \[x=\cos^{-1} \left( \frac{ 1 }{ \sqrt{5} } \right)\]

OpenStudy (anonymous):

To the nearest degree, that would be 63.4 degrees.

OpenStudy (firejay5):

the answer would be 63

OpenStudy (anonymous):

Ah, sorry, I'm used to rounding to the nearest tenth of a degree.

OpenStudy (firejay5):

it's tangent

OpenStudy (anonymous):